OpenStudy (anonymous):

How much water will fit in an aquarium that is shaped like a box and is 10 inches high, 6 inches wide, and 14 inches long?

OpenStudy (anonymous):

10*6*14 cubic inch = 840

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

l*b*h = volume of cuboid 10*6*14= 840 in (cubic)

OpenStudy (anonymous):

u are welcome and if u like my answer then give me a good answer too :d